OpenStudy (zmudz):

Let \(A = \begin{pmatrix}x_a\\y_a\\z_a\end{pmatrix}\), \(B= \begin{pmatrix}x_b\\y_b\\z_b\end{pmatrix}\), and \(C = \begin{pmatrix}x_c\\y_c\\z_c\end{pmatrix}\). Show that \((x_a,y_a,z_a), (x_b,y_b,z_b)\), and \((x_c,y_c,z_c)\) are collinear if and only if \(A \times B + B \times C + C\times A = 0.\) I'm totally stuck on this one. By the way, the \(0\) refers to the vector \(\begin{pmatrix} 0\\0\\0 \end{pmatrix}\)

OpenStudy (thomas5267):

By collinearity, \[ \mathbf{A}=\mathbf{d}+\lambda_1\mathbf{f}\\ \mathbf{B}=\mathbf{d}+\lambda_2\mathbf{f}\\ \mathbf{C}=\mathbf{d}+\lambda_3\mathbf{f}\\ \begin{align*} \mathbf{A}\times \mathbf{B}&=(\mathbf{d}+\lambda_1\mathbf{f})\times(\mathbf{d}+\lambda_2\mathbf{f})\\ &=\lambda_2\mathbf{d}\times\mathbf{f}+\lambda_1\mathbf{f}\times\mathbf{d}\\ \mathbf{B}\times \mathbf{C}&=\lambda_3\mathbf{d}\times\mathbf{f}+\lambda_2\mathbf{f}\times\mathbf{d}\\ \mathbf{C}\times \mathbf{A}&=\lambda_1\mathbf{d}\times\mathbf{f}+\lambda_3\mathbf{f}\times\mathbf{d} \end{align*} \] Add them together the first part follows. Let me think about the second part.